A 2.4 mol of ideal gas through an isochoric process in which the process in which the pressure changes from 1.8 atm to 2.9 atm at 200k. The specific heat of a gas at constant volume is 12.47j/mol.k . What was the initial temperature and the internal energy of the system?


1
Expert's answer
2021-06-17T15:08:02-0400
"T_2=T_1\frac{p_2}{p_1}\\
T_1=200\frac{1.8}{2.9}=124\ K"

"U_1=2.4(12.47)(124)=3711\ J"